OpenStudy (anonymous):

Rename 2/3 and 7/12 using the least common denominator.

OpenStudy (anonymous):

the denominators are... 3 and 12. what is the least common multple of 3 and 12?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

count by 3's count by 12's what's the first number common to both counts?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

that's the common denominator. so there's no need to do anything to the second fraction, it already has 12 as a denominator. to get the first with a denominator of 12, you need to multiply. \[\frac{ 2 }{ 3 }= \frac{ 2 }{ 3 }\cdot\left(1\right)= \frac{ 2 }{ 3 }\cdot\left(\frac{ 4 }{ 4 }\right)=\frac{ 8 }{ 12 }\] we can only multiply by 1 so we don't change the value, only the appearance. kind of like getting dressed up for halloween... you are still you, you just look different

OpenStudy (anonymous):

oh ok thank you!!